A 440 volt D.C. shunt motor has an armature resistance of 0.8 ohm and field resistance of 200 ohm. If motor is taking 20 Amp current from supply then the back e.m.f. will be
1
Expert's answer
2021-08-19T14:27:54-0400

Given Data:-

Supply Voltage V = 440 V

Line Current "I_L" = 20A

Shunt Resistance "R_{sh}" = 200Ω

Armature Resistance "R_a" = 0.8 Ω

Armature Current "I_a" =?

Back EMF "E_b" =?

Armature current

The Line current of DC shunt motor is the sum of Armature current and Shunt field current

"I_L=I_a+I_{sh}"

And Shunt Field Resistance is given as

"I_{sh}=\frac{V}{R_{sh}}=\frac{440}{200}=2.2A"

"I_a=I_L-I_{sh}=20-2.2=17.8A"


Back EMF

The Back EMF of DC Motor is given by

"E_b=V-I_aR_a=440-17.8\times 0.8=440-14.24=425.76V"